OpenStudy (anonymous):

Take the integral of (A/x)+(B/x^2)+(C/x^3)+(D/(x+1)^2)+(E/(x+1)^2)+(F/(x^2+8))

OpenStudy (anonymous):

might i suggest using wolfram integrator

OpenStudy (anonymous):

it didn't show me anything, can u help me

OpenStudy (anonymous):

you have to do it one a t a time first take the integral of A/x which is like A ln x then you do B /x^2 then C/x^3 etc etc etc this is assuming that ABCDEF are all constants and not an actual function

OpenStudy (anonymous):

do you not know the full thing?

OpenStudy (anonymous):

im lazy, i dont want to integrate A, B, and C are easy F you need to use trig substitution D and E look the same and you need to do a x+1=u substitution

OpenStudy (anonymous):

A/x = A ln(x)?

OpenStudy (anonymous):

Aln |x| thats something you should memorize

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{1}{x}dx= \ln |x|\]

OpenStudy (anonymous):

i don't even know what B would be since there is a ^2

OpenStudy (anonymous):

\[\frac{1}{x^2} = x^{-2}\]

OpenStudy (anonymous):

same thing goes for ^3

OpenStudy (anonymous):

@jim_thompson5910 @lgbasallote i want to sleep, please finish up

hero (hero):

^what he said